OpenStudy (anonymous):

y=107(4.5)^x type the model in terms of base e

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Recall that since the log and exponential function are inverses of one another:\[\Large\rm \color{orangered}{x}=e^{\ln \color{orangered}{x}}\] And so when we have an exponential function:\[\Large\rm \color{orangered}{a^x}\]We can write it as,\[\Large\rm e^{\ln(\color{orangered}{a^x})}\]So let's apply that to our problem and see what it looks like :)

OpenStudy (anonymous):

so what is my answer?

zepdrix (zepdrix):

So we apply the log and exponential composition to the problem: \[\Large\rm \color{orangered}{107(4.5)^x}=e^{\ln(\color{orangered}{107(4.5)^x})}\]And we need to apply some log rules from here so it looks a little better :)

OpenStudy (anonymous):

ok....

zepdrix (zepdrix):

Recall:\[\Large\rm \ln(a\cdot b)=\ln(a)+\ln(b)\]

zepdrix (zepdrix):

So then:\[\Large\rm \ln(107(4.5)^x)=\ln(107)+\ln(4.5^x)\]

OpenStudy (anonymous):

so is that my answer?

zepdrix (zepdrix):

No -_- why so impatient?

OpenStudy (anonymous):

because i really need to get this done please hurry

OpenStudy (anonymous):

now what?

OpenStudy (aum):

\[\large 4.5 = e^{\ln(4.5)} = e^{1.5041} \\ \large \text{ } \\ \large y=107(4.5)^x = 107(e^{1.5041})^x = 107e^{1.5041x} \]